OpenStudy (anonymous):

Find all the points of intersection of the graphs of f(x)=sinx+1 and g(x)=cosx on the interval {0,4pi}

OpenStudy (cwrw238):

solve for x: sinx + 1 = cos x

OpenStudy (jhannybean):

When you've done what @cwrw238 has instructed you with, you can check back on this graph :) http://fooplot.com/plot/f7e8c7awte

OpenStudy (anonymous):

you are right my friend @cwrw238 , but it's not such an easy equation to solve algebraically I recommend to set the equation in the form: cos(x)-sin(x)=1 this way we eliminate the first three quardrants because from the solution still we have the forth one, anyways cos(x)-sin(x) can never be 1 unless sin(x)=-1 , or cos(x)=1 this way we have the solutions :{0, 3pi/2 , 2pi , 7 pi/2 , 4pi}

OpenStudy (jhannybean):

I think he just wanted the person to work it out for themself xD

OpenStudy (anonymous):

$$\sin x+1=\cos x\\\sin^2 x+2\sin x+1=\cos^2 x\\\sin^2 x+2\sin x+1=1-\sin^2 x\\2\sin^2 x+2\sin x=0\\2(\sin x)(\sin x+1)=0$$now, when does \(\sin x=0\) or \(\sin x+1=0\)? make sure they satisfy the original equation!

OpenStudy (cwrw238):

i dont think thats too difficult an equation to solve. - although it depends of course on how much trigonometry avaob has done.

OpenStudy (anonymous):

you are right @cwrw238 ,, @oldrin.bataku has just solved it :)